# Exploring Self-Supervised Vision Transformers for Deepfake Detection: A Comparative Analysis

### v1

@@ Abstract
This paper investigates the effectiveness of self-supervised pre-trained transformers compared to supervised pre-trained transformers and conventional neural networks (ConvNets) for detecting various types of deepfakes. We focus on their potential for improved generalization, particularly when training data is limited. Despite the notable success of large vision-language models utilizing transformer architectures in various tasks, including zero-shot and few-shot learning, the deepfake detection community has still shown some reluctance to adopt pre-trained vision transformers (ViTs), especially large ones, as feature extractors. One concern is their perceived excessive capacity, which often demands extensive data, and the resulting suboptimal generalization when training or fine-tuning data is small or less diverse. This contrasts poorly with ConvNets, which have already established themselves as robust feature extractors. Additionally, training and optimizing transformers from scratch requires significant computational resources, making this accessible primarily to large companies and hindering broader investigation within the academic community. Recent advancements in using self-supervised learning (SSL) in transformers, such as DINO and its derivatives, have showcased significant adaptability across diverse vision tasks and possess explicit semantic segmentation capabilities. By leveraging DINO for deepfake detection with modest training data and implementing partial fine-tuning, we observe comparable adaptability to the task and the natural explainability of the detection result via the attention mechanism. Moreover, partial fine-tuning of transformers for deepfake detection offers a more resource-efficient alternative, requiring significantly fewer computational resources.
@@ 1 Introduction
In recent years, deepfake detection has emerged as a highly investigated field driven by the increasing prevalence of synthetic media [3]. Transfer learning, a commonly adopted strategy in computer vision, has also been widely used in deepfake detection [46, 37]. The selection of an appropriate backbone architecture plays an important role, serving not only as a feature extractor but also as a regularizer to prevent overfitting. Previous studies have predominantly relied on ConvNets that are pre-trained using supervised learning on ImageNet. However, with the recent advancements in transformer architectures [49], such as CLIP [35] and GPT-4 [1], particularly in multi-modal tasks, there has been growing interest in exploring their efficacy for deepfake detection. Despite their demonstrated success in various domains, the adoption of pre-trained ViTs [16] as feature extractors, especially large ones, has thus far been met with hesitation in the deepfake detection community. This reluctance stems from concerns about their immense capacity, which may exceed the requirements of the task and lead to potential overfitting, as well as their demanding resource requirements regarding training or fine-tuning data and computational resources. In contrast, ConvNets have already established themselves as robust feature extractors.
The advent of SSL has revolutionized the field of transformers, beginning with natural language processing (NLP) models such as BERT [25] and GPT [36]. Subsequently, DINO [5] showcased the successful adaptation of SSL on ViTs, resulting in robust feature extractors and enabling explicit semantic segmentation of images—an ability not readily available in supervised ViTs. The introduction of registers [13] in DINOv2 [34] further validated these capabilities, demonstrating their effectiveness in transfer learning across various downstream tasks. Specifically, in the realm of deepfake detection, the preliminary work of Cocchi et al. [11] demonstrated that detectors using either a basic k-NN classifier or a linear classifier equipped with pre-trained frozen DINO backbones can effectively identify images generated by Stable Diffusion models [38].
The current study substantially expands upon the findings of Cocchi et al. [11] in several key aspects. Our contributions can be summarized as follows:
  • We conduct an extensive comparative study to explore the utilization of pre-trained vision transformers in deepfake detection from two perspectives: utilizing their frozen backbones as multi-level feature extractors, a method increasingly utilized in the literature, and partially fine-tuning their final transformer blocks.
@@
  • We conclude that leveraging self-supervised learning on vision transformers, pre-trained using large datasets unrelated to deepfake detection, leads to a superior performance on the detection of various deepfake images and videos compared to utilizing supervised pre-training.
@@ 2.1 Self-supervised vision transformers
Caron et al. [5] argued that image-level supervision often oversimplifies rich visual information, reducing it to a single concept from a predefined category set. To address this limitation, they applied SSL on ViTs [16], particularly DeiT [47], to improve feature representation, leading to the development of DINO. This approach can be seen as a type of knowledge distillation [21] without explicit labels, leveraging techniques such as a momentum encoder [19], multi-crop training [4], and the utilization of small patches with ViTs. Through SSL training, DINO showcased a remarkable performance across various tasks, including image classification, image retrieval, copy detection, semantic layout discovery in scenes, video instance segmentation, probing the self-attention map, and transfer learning via fine-tuning on downstream tasks. Importantly, SSL provides explicit information about the semantic segmentation of the image, a capability not clearly present in supervised ViTs.
The introduction of DINOv2 [34] primarily focused on accelerating and stabilizing training at scale using a significantly larger dataset comprising images sourced from curated and uncurated data sources. A subsequent work by Darcet et al. [13] identified and characterized artifacts in the feature maps of supervised and self-supervised ViTs. They proposed a straightforward yet effective solution that augments the input sequence of ViTs with additional tokens, referred to as registers. These registers are utilized during training but discarded during inference. These enhancements contribute to the robustness and efficiency of training self-supervised ViTs, facilitating their broader application in various computer vision tasks, including deepfake detection.
@@ 2.2 Transformers in deepfake detection
In deepfake detection, transformers are primarily utilized in two ways: as feature refiners following ConvNets or as replacements for ConvNets as the main feature extractors.
The use of transformers as feature refiners has gained popularity following the introduction of the transformer architecture. In this approach, a ConvNet or an ensemble of ConvNets, often pre-trained and sometimes partially fine-tuned later, serves as the primary feature extractor. Subsequently, a transformer, typically a shallow one with few blocks, is trained to further refine the extracted features. Khan et al. [26] utilized XceptionNet [9] for feature extraction, followed by 12 transformer blocks for feature refining. Similarly, Wang et al. [50] utilized EfficientNet-B4 [44] as a feature extractor and introduced a multi-scale transformer as one branch while utilizing a frequency filter as another branch for additional processing of the extracted features. Coccomini et al. [12] opted for a smaller version of the feature extractor, EfficientNet-B0. Wang et al. [51] utilized ConvNets for feature extraction, followed by processing with a newly proposed convolutional pooling transformer before classification. Lin et al. [30] utilized a ConvNet for pre-processing, followed by a two-stream ConvNet and a transformer-based module. Notably, Zhao et al. [53] utilized XceptionNet as a feature extractor and proposed the interpretable spatial-temporal ViT, which comprises a decomposed spatial-temporal self-attention and a self-subtract mechanism to capture spatial artifacts and temporal inconsistency.
The second way of using transformers as the main feature extractors involves proposing novel architectures or leveraging the pre-trained large ViTs. In addition to combining patch embedding with EfficientNet-B7’s features for pre-processing, Heo et al. [20] applied a distillation method of DeiT [47] on a transformer for deepfake detection. Guan et al. [18] introduced a local sequence transformer that models temporal consistency on sequences of restricted spatial regions. Ojha et al. [33] implemented nearest neighbor and linear probing on the frozen supervised pre-trained CLIP’s ViT [35] for deepfake detection. Similarly, Cocchi et al. [11] utilized the same classifiers as Ojha et al. but additionally evaluated the self-supervised pre-trained DINO and DINOv2 as feature extractors. Liu et al. [31] developed a forgery-aware adapter integrated into a frozen CLIP’s ViT, adapting image features to discern and integrate local forgery traces within image and frequency domains. Motivated by the work of Ojha et al., Cocchi et al., and Liu et al. [33, 11, 31], we conduct a comparative analysis on the use of self-supervised ViTs for deepfake detection from two perspectives: 1) utilizing their frozen backbones as multi-level feature extractors and 2) partially fine-tuning their final transformer blocks.
Figure 1: Overview of the two investigated approaches.
@@ 3 Methodology
In this section, we introduce two approaches that leverage pre-trained ViTs, as illustrated in Fig. 1. Approach 1 generalizes previous methods in the literature (primarily applied to ConvNets) as a multi-level feature extractor on a frozen backbone. This approach has been widely adopted in the deepfake detection community [46, 37], often accompanied by sophisticated adaptors in recent work [31]. Notably, this strategy is endorsed as effective in the original DINO paper for various classification tasks [5], as well as in the detection of generative adversarial network (GAN) and diffusion images [33, 11]. Approach 2 involves fine-tuning the final transformer blocks, a technique less favored due to the perception that transformers have large capacities and a high number of parameters.
@@ 3.3 Approach 2: Fine-tuning last transformer blocks
This approach is more straightforward than Approach 1. We append the new classifier $\mathcal{C}$ after the backbone $\mathcal{B}$, as visualized on the right part of Fig. 1. This can be formalized by
@@
During fine-tuning, the first $n-k$ blocks are frozen. For the transformer backbones, the class (CLS) token and the register tokens (if they exist) are also unfrozen and fine-tuned along with the unfrozen $k$ final blocks and the new classifier $\mathcal{C}$.
@@ 4.1 Backbones
We meticulously selected ConvNet and transformer backbones widely utilized in the domains of computer vision and forensics, prioritizing models with comparable sizes to ensure equitable and informative comparisons. Detailed specifications are provided in Table 1.
In terms of backbone architectures, we opted for EfficientNetV2 [45], given its reputation as a robust ConvNet architecture and the popularity of its first version in the forensics community. Additionally, we selected two well-known transformer-based models: DeiT III, the supervised-learning variant of DINO, and EVA-CLIP [42], an enhanced version of the renowned CLIP [35]. For supervised learning, we incorporated two prevalent annotations: the conventional class labels annotation and the multimodal image-text pairs annotation. We utilized the official pre-trained weights provided by the authors.
@@ 4.2 Datasets
We followed the data design of Nguyen et al. [32] by gathering a variety of images generated or manipulated by various deepfake methods to construct the main datasets. The details of the training, validation, and test sets are shown in Table 2. The datasets were designed to be balanced regarding the ratio of real and fake images and the number of images per training method, and were guaranteed not to overlap.
Table 2: Sizes of the main training, validation (val), and test (seen) sets, inspired by Nguyen et al.’s design [32], along with the sizes of the unseen validation and test sets from Țânțaru et al.’s dataset [55].
@@
Real images were gathered from the VidTIMIT [40], VoxCeleb2 [10], FaceForensics++ (FF++) [39], Google DFD [17], Deepfake Detection Challenge Dataset (DFDC) [15], and Celeb-DF [29] datasets.
One part of the fake images comprised images gathered from the FF++, Google DFD, Celeb-DF, DFDC, DeepfakeTIMIT (DF-TIMIT) [27], and YouTube-DF [28] datasets. The other part of the fake images were images generated by various GANs, including StarGAN [7], StarGAN-v2 [8], RelGAN [52], ProGAN [22], StyleGAN [23], and StyleGAN2 [24].
@@
4.3 Metrics
We used five metrics in our evaluation:
  • Classification accuracy $=\tfrac{{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$, where TP, TN, FP, and FN are true positive, true negative, false positive, and false negative, respectively.
  • True positive rate (TPR) $=\tfrac{TP}{TP+FN}$.
  • True negative rate (TNR) $=\tfrac{TN}{TN+FP}$.
  • Equal error rate (EER): value when false positive rate (FPR) equals false negative rate (FNR).
  • Half total error rate (HTER): $\text{HTER}=\tfrac{\text{FPR}+\text{FNR}}{2}$.
5 Results and discussion
In this section, we first discuss Approaches 1 and 2 in Sections 5.1 and 5.2, respectively. We compare the performances among different architectures and versions of DINO, as well as between DINOs and ConvNets and other transformers in more detail. Additionally, we implement improvements and conduct ablation studies on the best performing architectures to gain further enhancements and insights. Next, we evaluate selected models on the unseen test set in Sec. 5.3. Lastly, we visualize the attention maps of the fine-tuned DINO model and its original SSL pre-trained version on real and deepfake image examples in Sec. 5.4.
We initially validated the findings of Cocchi et al. [11] with some minor extensions by applying conventional classifiers to the features extracted by frozen SSL pre-trained DINO backbones (with small size (S) and base size (B)), specifically extracted from the CLS token. In addition to the nearest neighbors (k-NN) classifier and linear probing, we utilized k-means on the principal component analysis (PCA) features [32] and a two-layer perceptron classifier.
@@ 5.1 Approach 1: Using frozen backbone as a multi-level feature extractor
Deepfake detection involves the identification of deepfake fingerprints, such as artifacts or irregular patterns. Therefore, relying solely on the CLS token may not be optimal. We therefore assessed the effectiveness of incorporating patch tokens and multiple intermediate features extracted from the final $k$ blocks instead of solely from the final block. Additionally, we compared the performance of two feature fusion techniques: weighted sum (WS) and concatenation (concat). Since utilizing all tokens (CLS and patch tokens) results in huge feature sizes, we only evaluated the weighted sum technique in this case. Regarding the DINO backbones, in addition to small and base sizes, we also evaluated large (L) and giant (G) sizes, which are only publicly available in DINOv2. The results are presented in Table 4.
@@
We selected the DINOv2 - ViT-L/14-Reg (chosen for its balance between performance and model size) to assess potential enhancements. Simple linear adaptors were utilized to reduce feature dimensionality and enable feature concatenation. Additionally, dropout was applied to mitigate overfitting. The results are presented in Table 5. The optimal configuration involves using dropout alongside linear adaptors and feature concatenation.
Table 6: Comparison between different ConvNet and transformer architectures on Approach 1. The final setting in Table 5 was applied. Accuracies were calculated using a threshold of 0.5.
Architecture Accuracy EER HTER
EfficientNetV2 Large 78.67 21.77 21.45
DeiT III L/16-LayerScale 79.96 19.77 19.96
EVA-02-CLIP-L/14 83.31 16.51 16.76
DINOv2 ViT-L/14-Reg 87.42 11.98 12.41
We applied the optimal configuration to EfficientNetV2, DeiT III, and EVA-CLIP and then compared their performances with that of DINOv2. The results are displayed in Table 6. DINOv2 clearly outperformed EfficientNetV2 and DeiT III, and it surpassed EVA-CLIP despite the latter’s pre-training on a larger dataset with rich annotations (image-text pairs). These results underscore the advantage of using SSL for pre-training, enabling the learning of superior representations applicable to multiple tasks.

### v2

@@ Abstract
This paper investigates the effectiveness of self-supervised pre-trained vision transformers (ViTs) compared to supervised pre-trained ViTs and conventional neural networks (ConvNets) for detecting facial deepfake images and videos. It examines their potential for improved generalization and explainability, especially with limited training data. Despite the success of transformer architectures in various tasks, the deepfake detection community is hesitant to use large ViTs as feature extractors due to their perceived need for extensive data and suboptimal generalization with small datasets. This contrasts with ConvNets, which are already established as robust feature extractors. Additionally, training ViTs from scratch requires significant resources, limiting their use to large companies. Recent advancements in self-supervised learning (SSL) for ViTs, like masked autoencoders and DINOs, show adaptability across diverse tasks and semantic segmentation capabilities. By leveraging SSL ViTs for deepfake detection with modest data and partial fine-tuning, we find comparable adaptability to deepfake detection and explainability via the attention mechanism. Moreover, partial fine-tuning of ViTs is a resource-efficient option.
@@ 1 Introduction
In recent years, facial deepfake detection has emerged as a highly investigated field driven by the increasing prevalence of synthetic media [3]. Transfer learning, a commonly adopted strategy in computer vision, has also been widely used in deepfake detection [48, 39]. The selection of an appropriate backbone architecture plays an important role, serving not only as a feature extractor but also as a regularizer to prevent overfitting. Previous studies have predominantly relied on ConvNets that are pre-trained using supervised learning on ImageNet. However, with the recent advancements in transformer architectures [51], such as CLIP [37] and GPT-4 [1], particularly in multi-modal tasks, there has been growing interest in exploring their efficacy for deepfake detection. Despite their demonstrated success in various domains, the adoption of pre-trained ViTs [17] as feature extractors, especially large ones, has thus far been met with hesitation in the deepfake detection community. This reluctance stems from concerns about their immense capacity, which may exceed the requirements of the task and lead to potential overfitting, as well as their demanding resource requirements regarding training or fine-tuning data and computational resources. In contrast, ConvNets have already established themselves as robust feature extractors.
The advent of SSL has revolutionized the field of transformers, beginning with natural language processing (NLP) models such as BERT [27] and GPT [38]. Subsequently, DINO [5] and masked autoencoders (MAEs) [20] showcased the successful adaptation of SSL on ViTs, resulting in robust feature extractors and enabling explicit semantic segmentation of images—an ability not readily available in supervised ViTs. The introduction of registers [13] in DINOv2 [36] further validated these capabilities, demonstrating their effectiveness in transfer learning across various downstream tasks. Specifically, in the realm of deepfake detection, the preliminary work of Cocchi et al. [11] demonstrated that detectors using either a basic k-NN classifier or a linear classifier equipped with pre-trained frozen DINO backbones can effectively identify images generated by Stable Diffusion models [40].
The current study substantially expands upon the findings of Cocchi et al. [11] in many key aspects. With the focused on comparing various backbones as feature extractors, i.e., evaluating the quality of their extracted representations for deepfake detection, our contributions can be summarized as follows:
  • We conducted a comparative study on using pre-trained ViTs in facial deepfake detection from two perspectives: using their frozen backbones as multi-level feature extractors and partially fine-tuning their final transformer blocks. We employed simple classifiers as a proof-of-concept, which offers two benefits: reducing nuisance factors that could influence the comparison and ensuring that the results can generalize to any downstream classifier, whether simple or complex.
@@
  • We conclude that leveraging SSL on ViTs, particularly DINOs, pre-trained using large datasets unrelated to deepfake detection, leads to a superior performance on the detection of various deepfakes compared to utilizing supervised pre-training.
@@ 2.1 Self-supervised vision transformers
Caron et al. [5] argued that image-level supervision oversimplifies rich visual information to a single concept. They applied SSL on a ViT [17] called DeiT [49] to improve feature representation, resulting in DINO. This approach can be seen as a type of knowledge distillation [23] without explicit labels, leveraging techniques such as a momentum encoder [21], multi-crop training [4], and small patches with ViTs. Subsequently, He et al. used a masking mechanism for training MAEs, similar to BERT’s [27]. SSL training with DINO and MAE showcased remarkable performance across various vision tasks and transfer learning via fine-tuning on downstream tasks. Importantly, DINO offered explicit information about semantic segmentation, a capability not clearly present in supervised ViTs.
The introduction of DINOv2 [36] focused on accelerating and stabilizing training at scale using a larger dataset from curated and uncurated sources. Darcet et al. [13] identified artifacts in the feature maps of supervised and self-supervised ViTs and proposed a solution that augments the input sequence with additional tokens, called registers. These registers are used during training but discarded during inference. These enhancements improve the robustness and efficiency of training self-supervised DINO, facilitating their application in various computer vision tasks, including deepfake detection.
@@ 2.2 Transformers in deepfake detection
In deepfake detection, transformers are primarily used in two ways: as feature refiners following ConvNets or as replacements for ConvNets as the main feature extractors.
The use of transformers as feature refiners has gained popularity since the introduction of the transformer architecture. In this approach, a ConvNet or an ensemble of ConvNets, often pre-trained and sometimes partially fine-tuned, serves as the primary feature extractor. A transformer, typically shallow with few blocks, is then trained to refine the extracted features. Khan et al. [28] used XceptionNet [9] for feature extraction, followed by 12 transformer blocks for feature refining. Similarly, Wang et al. [52] used EfficientNet-B4 [46] as a feature extractor and introduced a multi-scale transformer as one branch while utilizing a frequency filter as another branch for additional processing of the extracted features. Coccomini et al. [12] opted for the smaller extractor (EfficientNet-B0). Wang et al. [53] used ConvNets followed by a proposed convolutional pooling transformer before classification. Lin et al. [32] used a ConvNet for pre-processing, followed by a two-stream ConvNet and a transformer-based module. Notably, Zhao et al. [55] used XceptionNet as a feature extractor and proposed the interpretable spatial-temporal ViT with decomposed spatial-temporal self-attention and a self-subtract mechanism to capture spatial artifacts and temporal inconsistency.
The second way of using transformers as the main feature extractors involves proposing novel architectures or leveraging the pre-trained large ViTs. Heo et al. [22] combined patch embedding with EfficientNet-B7’s features for pre-processing and applied a distillation method of DeiT [49] on a transformer for deepfake detection. Guan et al. [19] introduced a local sequence transformer to model temporal consistency on sequences of restricted spatial regions. Ojha et al. [35] implemented nearest neighbor and linear probing on the frozen supervised pre-trained CLIP’s ViT [37] for deepfake detection. Cocchi et al. [11] used the same classifiers as Ojha et al. but additionally evaluated the self-supervised pre-trained DINO and DINOv2 as feature extractors. Liu et al. [33] developed a forgery-aware adapter integrated into a frozen CLIP’s ViT, adapting image features to discern and integrate local forgery traces within image and frequency domains. Das et al. [14] borrowed the concept of MAE for the self-supervised auxiliary task. Motivated by the work of Ojha et al., Cocchi et al., and Liu et al. [35, 11, 33], we conduct a comparative analysis on the use of self-supervised ViTs for deepfake detection from two perspectives: 1) utilizing their frozen backbones as multi-level feature extractors and 2) partially fine-tuning their final transformer blocks.
Figure 1: Overview of the two investigated approaches. Blue blocks mean frozen blocks, while orange blocks mean fine-tuned or trained blocks.
@@ 3 Methodology
We introduce two approaches leveraging pre-trained ViTs, as illustrated in Fig. 1. Approach 1 is a generalized form that uses a frozen backbone as a feature extractor, widely adopted in deepfake detection [48, 39] and often paired with sophisticated adaptors [33]. This strategy is endorsed in the original DINO and MAE papers for various classification tasks [5, 20] and in detecting generative adversarial network (GAN) and diffusion images [35, 11]. Approach 2 involves fine-tuning the final transformer blocks, a less common technique due to the perception of transformers having large capacities and many parameters.
@@ 3.1 Problem formalization
DINO (various versions) [5] Transformer Self-supervised ImageNet-21K [15] 14M Not used
@@ 3.3 Approach 2: Fine-tuning last transformer blocks
This approach is more straightforward than Approach 1. We append the new classifier $\mathcal{C}$ after the backbone $\mathcal{B}$, as visualized on Fig. 1’s right part. This can be formalized by
@@
During fine-tuning, the first $n-k$ blocks are frozen. The class (CLS) token and register tokens (if exist) are also fine-tuned along with the last $k$ blocks and the new classifier $\mathcal{C}$.
@@ 4.1 Backbones
We meticulously selected ConvNet and transformer backbones widely utilized in the domains of computer vision and forensics, prioritizing models with comparable sizes to ensure equitable and informative comparisons. Detailed specifications are provided in Tab. 1.
For ConvNets’ representative, we opted for EfficientNetV2 [47] due to its robust architecture and the popularity of its predecessor in the forensics community. For supervised ViTs, we selected two well-known models: DeiT III [50] (having the same architecture as DINO, using the conventional class labels annotation) and EVA-CLIP [44] (an enhanced version of the renowned CLIP [37], using multimodal image-text pairs annotation). For SSL ViT, we chose DINOs [5, 36, 13] and MAE [20], focusing more on DINOs for simplicity. We utilized the official pre-trained weights provided by the authors.
@@ 4.2 Datasets
We followed the data design of Nguyen et al. [34] by gathering a variety of images generated or manipulated by various deepfake methods to construct the main datasets. The details of the training, validation, and test sets are shown in Tab. 2. The datasets were designed to be balanced regarding the ratio of real and fake images and the number of images per training method, and were guaranteed not to overlap.
Table 2: Sizes of the main training, validation (val), and test (seen) sets, inspired by Nguyen et al. [34], and of the unseen validation and test sets from Țânțaru et al. [57].
@@
Real images were gathered from the VidTIMIT [42], VoxCeleb2 [10], FaceForensics++ (FF++) [41], Google DFD [18], Deepfake Detection Challenge Dataset (DFDC) [16], and Celeb-DF [31] datasets. One part of the fake images comprised images gathered from the FF++, Google DFD, Celeb-DF, DFDC, DeepfakeTIMIT (DF-TIMIT) [29], and YouTube-DF (YT-DF) [30] datasets. The other part were images generated by various GANs, including StarGAN [7], StarGAN-v2 [8], RelGAN [54], ProGAN [24], StyleGAN [25], and StyleGAN2 [26].
@@
4.3 Metrics
We used some or all of the following five to measure the performance of the detectors: classification accuracy, true positive rate (TPR), true negative rate (TNR), equal error rate (EER), and half total error rate (HTER).
@@ 4.3 Metrics
5 Results and discussion
In this section, we first discuss Approaches 1 and 2 in Sections 5.1 and 5.2, respectively. We compare the performances among different architectures and versions of DINO, as well as between DINOs versus ConvNets and other ViTs. Additionally, we implement improvements and conduct ablation studies on the best-performing architectures to gain further enhancements and insights. Next, we evaluate selected models on the unseen test set in Sec. 5.3. Lastly, we visualize the attention maps of the fine-tuned DINOv2 model versus their original SSL pre-trained versions on real and deepfake image examples in Sec. 5.4.
We initially validated the findings of Cocchi et al. [11] with some minor extensions by applying conventional classifiers to the features extracted by frozen SSL pre-trained DINO backbones (with small size (S) and base size (B)), specifically extracted from the CLS token. In addition to the nearest neighbors (k-NN) classifier and linear probing, we utilized k-means on the principal component analysis (PCA) features [34] and a two-layer perceptron classifier. The results are shown in Tab. 3.
Deepfake detection involves identifying deepfake fingerprints, such as artifacts or irregular patterns. Therefore, relying solely on the CLS token may not be optimal. We assessed the effectiveness of incorporating patch tokens and multiple intermediate features from the final $k$ blocks rather than solely from the final block. We also compared two feature fusion techniques: weighted sum (WS) and concatenation (concat). Given the large feature sizes when using all tokens (CLS and patch tokens), we only evaluated the weighted sum technique in this case. For the DINO backbones, we evaluated small (S), base (B), large (L), and giant (G) sizes, the latter two being available only in DINOv2. The results are presented in Tab. 4.
@@ 5.1 Approach 1: Using frozen backbone as a multi-level feature extractor
Table 6: Comparison between different ConvNet and transformer architectures on the seen test set. For Approach 1, the final setting in Tab. 5 was applied. $k$ denotes the number of final blocks utilized for feature extraction in Approach 1 while signifies the number of fine-tuned blocks in Approach 2. Accuracies were calculated using a threshold of 0.5.
[…]
FF++
Real
[…]
FF++
DF
[…]
FF++
F2F
[…]
FF++
FS
[…]
FF++
NT
[…]
FF++
FSh
[…]
DFD
Real
[…]
DFD
Fake
[…]
Vid-
TIMIT
[…]
DF-
TIMIT
[…]
Vox-
Celeb2
[…]
YT-
DF
[…]
DFDC
Real
[…]
DFDC
Fake
[…]
Approach 1
EfficientNetV2 Large 4 60.90 84.70 83.40 79.50 75.30 83.50 61.00 85.50 59.75 67.90 92.62 67.40 54.00 71.60 95.55 78.50 21.71 21.62
DeiT III L/16-LayerScale 4 38.80 88.00 87.10 81.60 82.20 82.70 37.60 92.00 93.45 39.90 99.68 53.10 57.20 62.20 95.70 79.96 19.77 19.96
EVA-02-CLIP-L/14 4 50.40 97.40 90.80 93.20 80.80 90.30 44.50 97.20 83.30 66.40 99.82 31.90 57.10 83.90 99.90 83.30 16.51 16.77
[…]
DINOv2 ViT-L/14-Reg 4 72.40 94.60 87.00 89.30 76.50 89.40 69.40 98.30 96.80 46.20 99.70 67.30 77.70 76.00 99.85 87.42 11.98 12.41
Approach 2
EfficientNetV2 Large 1 65.00 94.60 89.00 89.30 84.10 90.50 22.90 91.50 99.20 60.40 99.16 62.50 72.40 68.90 97.25 84.75 15.05 15.22
DeiT III L/16-LayerScale 1 56.90 97.00 89.50 89.90 84.10 86.50 33.40 95.00 96.35 18.60 99.08 56.50 66.40 79.00 97.25 82.64 17.21 17.28
EVA-02-CLIP-L/14 1 53.50 97.50 92.70 92.40 83.30 92.40 62.40 98.80 96.55 62.90 99.26 82.20 64.00 84.60 99.00 88.29 11.77 11.77
[…]
DINOv2 ViT-L/14-Reg 1 75.60 97.20 92.80 94.80 81.60 93.40 30.80 99.60 99.75 62.70 99.74 76.90 74.90 86.10 99.85 88.78 11.32 11.26
[…]
DINOv2 ViT-L/14-Reg 11 85.10 98.60 94.30 95.40 89.70 97.20 67.50 99.60 99.85 92.40 99.86 86.80 88.10 89.00 99.65 94.38 5.63 5.64
We selected the DINOv2 - ViT-L/14-Reg (chosen for its balance between performance and model size) to assess potential enhancements. Simple linear adaptors were utilized to reduce feature dimensionality and enable feature concatenation. Additionally, dropout was applied to mitigate overfitting. The results are presented in Tab. 5. The optimal configuration involves using dropout alongside linear adaptors and feature concatenation.
We applied the optimal configuration on the selected backbones and then compared their performances. The results are displayed in Tab. 6. DINOv2 and MAE clearly outperformed EfficientNetV2 and DeiT III, and it surpassed EVA-CLIP despite the latter’s pre-training on a larger dataset with rich annotations (image-text pairs). Between the two SSL models, MAE was slightly better than DINOv2. These results underscore the advantage of using SSL for pre-training, enabling the learning of superior representations applicable to multiple tasks.